OpenStudy (anonymous):

Help with rules of integration t*(1.1)^t dt

zepdrix (zepdrix):

More Parts! :)

zepdrix (zepdrix):

\[\large\rm u=t,\qquad\qquad dv=\left(1.1\right)^t~dt\]Do you remember how to find your v? How to integrate this exponential?

OpenStudy (anonymous):

ummm its ln 1.1 - ummm ughhh no i don't remember.

zepdrix (zepdrix):

General rule for differentiating exponentials:\[\large\rm \frac{d}{dx}a^x\quad=\quad a^x \ln(a)\]

zepdrix (zepdrix):

We multiply by the log of the base. So when we go the other direction, integration, \[\large\rm \int\limits a^x~dx=\frac{1}{\ln(a)}a^x\]We divide by the log of the base.

OpenStudy (anonymous):

so then we have t(1.1)^t - integral (1.1^t)/ln (1.1)

OpenStudy (anonymous):

ohhh no thats not it

zepdrix (zepdrix):

Let's get our 4 pieces gathered together so there is no confusion,\[\large\rm u=t,\qquad\qquad dv=\left(1.1\right)^t~dt\]\[\large\rm du=dt,\qquad\qquad v=\frac{(1.1)^t}{\ln(1.1)}\]

zepdrix (zepdrix):

And I guess you're trying to set up:\[\large\rm uv-\int\limits v~ du\]ya?

OpenStudy (anonymous):

ya... so now the integral of v. which is ... gimme a sec

OpenStudy (anonymous):

(1.1)^t/ln(1.1)^2

zepdrix (zepdrix):

Good good good, you just get another one of those denominators!

OpenStudy (anonymous):